# Your Paper Title Here

### v1

@@ Abstract
Machine learning models trained on Earth observation data, such as satellite imagery, have demonstrated significant promise in predicting household-level wealth indices, enabling the creation of high-resolution wealth maps that can be leveraged across multiple causal trials while addressing chronic data scarcity in global development research. However, because standard training objectives prioritize overall predictive accuracy, these predictions inherently suffer from shrinkage toward the mean, leading to attenuated estimates of causal treatment effects and limiting their utility in policy evaluations. Existing debiasing methods, such as Prediction-Powered Inference (PPI), can handle this attenuation bias but require additional fresh ground-truth data at the downstream stage of causal inference, which restricts their applicability in data-scarce environments. In this paper, we introduce and evaluate two correction methods—linear calibration correction and Tweedie’s correction—that substantially reduce prediction bias without relying on newly collected labeled data. Linear calibration (LCC) corrects bias through a straightforward linear transformation derived from held-out calibration data, whereas Tweedie’s correction leverages empirical Bayes principles to directly address shrinkage-induced biases by exploiting score functions derived from evaluating the model’s learning patterns. Through analytical exercises and experiments using Demographic and Health Survey (DHS) data, we demonstrate that both proposed methods meet or outperform existing approaches that either require (a) adjustments to training pipelines or (b) additional labeled data, achieving significant reductions in attenuation bias in data-scarce environments. These approaches may represent a promising avenue for improving the reliability of causal inference when direct outcome measures are limited or unavailable, enabling a “one map, many trials” paradigm where a single upstream data creation team produces predictions usable by many downstream teams across diverse ML pipelines.
@@ Introduction
Machine learning (ML) models trained on Earth observation (EO) data have recently demonstrated impressive performance in predicting household-level wealth indices, such as the International Wealth Index (IWI), with reported $R^{2}$ values approaching 0.80 (Pettersson et al., 2023; Jean et al., 2016; Daoud et al., 2023; Chi et al., 2022). These EO-ML models offer a much-needed solution to longstanding data scarcity issues in global development, enabling researchers to generate high-resolution, retrospective “wealth maps” in data-poor regions. But while these predictions are promising in the EO-ML upstream task, a key question remains (Jerzak et al., 2023a, b): the value of EO-ML-derived data hinges on whether domain researchers, in the downstream phase, can use these data to estimate causal effects for their specific application area downstream of the research process, from evaluating applied environmental to social science questions Daoud and Dubhashi (2023).
However, even under ideal conditions, the predictions of EO-ML models tend to have a smaller variance than the target values, leading to attenuated treatment effects and, thus, potentially misleading conclusions (Berglund, 2012).
@@
Recent research has developed methods that can debias the upstream EO-ML model, but they require either fresh data in the downstream Olofsson et al. (2013); Egami et al. (2023); Lu et al. (2024) or changing the way the model is trained Ratledge et al. (2022). For example, relying on such fresh data, Angelopoulos et al. (2023) proposes Prediction-Powered Inference (PPI), a method that corrects any ML model for downstream inference. Similarly, Egami et al. (2023) achieves the same goal but with better sample efficiency de Pieuchon et al. (2025), but it also requires knowing the probability of observing a sample point. While these methods offer a path to integrate upstream and downstream analyses, their success depends on the sometimes availability of newly collected treatment-predictor-outcome tuples in the downstream causal step or knowledge of the data labeling process (Egami et al., 2023). Often, such data are prohibitively costly or unethical to collect.
@@
Ratledge et al. (2022) addresses the challenge of attenuation bias and the unavailability of data in the downstream by adjusting the model’s loss function instead. During training, they penalize quintile-level discrepancies between observed and predicted means, encouraging predictions to match the true outcome distribution more closely across its support. Although their solutions do not need additional downstream data, they require modifying the training pipeline, which is likely computationally expensive.
@@
Thus, to enable broader applicability of ML models, the upstream team needs to create a data product that is agnostic to downstream use, requires no downstream-fresh data, produces unbiased causal estimates (due to attenuation), and is computationally efficient Daoud and Dubhashi (2023).
@@
We address this gap by proposing Tweedie’s correction, a novel method that reduces prediction bias without requiring fresh labeled data. We evaluate this approach against four others: linear calibration correction, Prediction-Powered Inference (PPI), the loss adjustment approach of Ratledge et al. (2022), and naive uncorrected predictions. The Tweedie’s correction paves the way for the same wealth map to be reused across multiple trials without needing to refit a model for each trial. This also removes the requirement for interaction between the upstream machine-learning team creating the data product and any other domain-knowledge team in the downstream phase. Hence the concept: one map, many trials. We derive analytical proofs and demonstrate the empirical performance of all methods in both simulations and real-world applications, using data from the African Demographic and Health Survey (DHS).
Figure 1: Illustrating the proposed debiasing pipeline for causal inference using machine learning predictions from satellite imagery, without requiring fresh ground-truth data.
1 Problem Setup & Related Work
@@
Statistical predictions can systematically shrink estimates toward the central modes of the training distribution (Ting, 2024a). This phenomenon is rooted in optimal error-minimization estimation (cf. Stein’s paradox), where a model “hedges” predictions in regions of greater uncertainty. When facing noise or limited training data, the optimal strategy, in terms of minimizing residuals, is to pull predictions toward denser regions of the training data. For instance, consider an image that appears with an IWI of 80, but the model has observed this outcome only once, compared to thousands of observations with an IWI of around 70. A perfect prediction would be 80 for this data point, but the model will pull this data point’s prediction towards 70 instead to hedge against the risk of overcommitting to a rare value.
@@
When such biased predictions replace true outcomes in downstream causal analyses, the resulting regression-to-the-mean induces attenuation bias, reducing observed treatment group differences even under randomization (Shu and Yi, 2019). The difference in mean predicted outcomes between treated and control units shrinks both group means toward the global average, thereby attenuating the estimated effect.
@@
Regression-to-the-mean bias is particularly acute in upstream tasks with scarce ground truth (due to finite samples) or in high-dimensional input data (due to the curse of dimensionality), forcing greater reliance on error-prone predictions—something especially pertinent in development economics, political science, and global health contexts —where household data is often expensive or unavailable—to increasingly substitute upstream predictions (Kino et al., 2021; Kakooei et al., 2024; Daoud and Johansson, 2024). Moreover, high out-of-sample $R^{2}$ does not eliminate this bias: models can achieve strong accuracy yet yield attenuated causal estimates (Ting, 2024b).
@@
Existing remedies, such as PPI, which subtracts a rectifier estimated from fresh ground-truth pairs at the causal-inference stage (Angelopoulos et al., 2023), or loss-function adjustments that penalize quintile-level errors (Ratledge et al., 2022). However, these approaches require that the upstream team adapts the data for downstream analysis, slowing or limiting altogether "one map, many trials" efforts.
@@
To address these limitations, we propose a Tweedie-inspired method. This Tweedie method debiases the shrinkage effect in ML-predicted outcomes, enabling downstream researchers to accurately recover causal effects from the upstream data product. Our Tweedie approach treats the ML model as a black box: we assume access only to its out-of-sample predictions and do not rely on retraining or modifying the model. Under this setting, we demonstrate that shrinkage can be reversed, either by learning a global calibration correction from held-out labeled data or by estimating a local correction using ideas from Tweedie’s formula, initially described by Robbins and Tweedie in 1956. Both strategies aim to recover unbiased treatment effect estimates from biased predictions, without requiring fresh ground-truth labels at the time of intervention. We benchmark our Tweedie adaptation to existing methods, which are compared in Table A.III.1 based on their associated assumptions and data/training requirements. As previously mentioned, the only method we are unable to deploy in our setting is the DSL method, as the probability of DHS samples is a function of unknown aid disbursement strategies.
Setup
@@
Our pipeline assumes access to two datasets. First, in the upstream, a labeled dataset with training points (with indices in $\mathcal{I}_{\textrm{Train}}$): $D_{\text{Train}}=\{(\bm{X}_{i},Y_{i}):i=\mathcal{I}_{\textrm{Train}}\},$ where $\bm{X}_{i}\in\mathcal{X}$ denotes covariates (e.g., satellite imagery) and $Y_{i}$ is the true, ground-truth outcome (e.g., IWI poverty level as from the DHS).
@@
Next, in the downstream phase, we assume that different teams have assembled collections of trial datasets indexed by $s$, corresponding to different randomized controlled trials, quasi-experiments, or geographic partitions, with a specific treatment (or action) of interest indexed by treatment setting: $D_{\text{Trial $s$}}=\{(\bm{X}_{i},\widehat{Y}_{i},A_{is}):i\in\mathcal{I}_{\textrm{Trial $s$}},s\in\mathcal{S}\}.$ The data product associated with $Y_{i}$ is general (e.g., all-Africa), but investigators may want to use that data in multiple causal settings, $s$, each with a different treatment, $A_{is}\in\{0,1\}$. The outcome $Y_{i}$ is not directly observed for all $i$. Instead, inference will need to take place through the noisy variable, $\hat{Y}_{i}$. We henceforth remove dependence on $s$ for simplicity.
@@
A key idea henceforth is that the upstream team—the one machine-learning team—creates an agonstic dataset (Grimmer et al., 2022), meaning that this team’s data product should produce unbiased causal estimates for any number of applications (trials) for the variety of teams downstream. This implies that there is no communication between up and downstream teams–hence the firewall between the two phases.
@@ 2 Methodology
In this section, we present a suite of methods designed to debias treatment effect estimates that rely on ML-predicted outcomes. Our overarching goal is to mitigate regression-to-the-mean dynamics in the downstream analysis, improving estimates of causal effects or population-level statistics. We first more formally describe two existing solutions, PPI and the method of Ratledge et al. (2022), before introducing two no-new-data debiasing approaches—Linear Calibration Correction (LCC) and Tweedie’s Correction—that use held-out data (without requiring fresh labels in the causal step) to correct shrinkage bias.
@@
PPI proposes a framework for valid statistical inference when a large set of unlabeled observations is paired with a predictive model’s outputs (Angelopoulos et al., 2023). Concretely, letting $\theta^{*}$ be the target parameter (e.g., a population mean) and $f$ denote the model that maps covariates $\bm{X}_{i}$ to predicted outcomes $f(\bm{X}_{i})$, PPI forms an estimator by adding a rectifier term—an adjustment for the average discrepancy between $Y_{i}$ and $f(\bm{X}_{i})$ observed in a smaller, labeled dataset. An estimate for $\theta^{*}$ is computed using an “imputed” version based on unlabeled data ($\{\bm{X}_{i}\}_{i={n+1}}^{N}$), then corrected by subtracting the bias estimated on the labeled subset. For mean estimates, this adjustment is:
$\hat{\theta}^{\textrm{PPI}}=\frac{1}{N-n}\sum_{i={n+1}}^{N}f(\bm{X}_{i})-\frac{1}{n}\sum_{i=1}^{n}\left(f(\bm{X}_{i})-Y_{i}\right)$
@@
Because the rectifier is estimated on ground-truth pairs $\{\bm{X}_{i},Y_{i}\}_{i=1}^{n}$, the procedure maintains valid coverage for $\theta^{*}$ in large samples, provided that the predictive model $f$ is fixed (i.e., not retrained on the labeled data). Unlike our approach, however, PPI crucially relies on gathering labeled observations in the downstream phase. Such labels come from high-quality surveys coordinated by USAID across the African continent and cost millions of US dollars to execute. Given the recent funding cuts to USAID, it is unlikely that new and fresh DHS data points will be collected. Hence, there is an acute need for unbiased data collection, which should be fully handled in the upstream phase. We will nonetheless include PPI correction in our analysis with the goal of showing what would have happened to our estimates had fresh data actually been available downstream.
@@
Another approach to remedy attenuation bias is to adjust the loss function: Ratledge et al. (2022) propose a specialized upstream training procedure that directly penalizes quintile-level discrepancies between observed and predicted means, thereby encouraging $\hat{Y}_{i}$ to match the true distribution of $Y_{i}$ more closely across the support of the outcome distribution. Specifically, this approach modifies the model’s objective function, punishing a quintile having a higher MSE than the others:
$\mathcal{L}=\mathrm{MSE}+\lambda_{b}\hat{E}_{b},$
@@
MSE is the mean squared error, $\lambda_{r}L_{2}$ is the L2 regularization term, and $\lambda_{b}\hat{E}_{b}$ is the quintile-specific bias loss, with $\hat{E}_{b}=\max_{j}(\hat{B}_{j}(\hat{f})^{2})$ and $\hat{B}_{j}(\hat{f})=\mathbb{E}[\hat{f}(\bm{X}_{i})-Y_{i}\mid Y_{i}\in Q_{j}]$. By design, this approach reduces the likelihood that group-level estimates of causal effects are attenuated when $\hat{Y}_{i}$ is used in downstream analyses. In contrast, our methods operate as post-hoc corrections on out-of-sample predictions from a black-box model, requiring no specialized retraining or customized loss function that might reduce overall predictive performance.
2.1 Linear Calibration Correction (LCC)

### v2

# Debiasing Machine Learning Predictions for Causal Inference Without Additional Ground Truth Data: “One Map, Many Trials” in Satellite-Driven Poverty Analysis

@@ Abstract
Machine learning models trained on Earth observation data, such as satellite imagery, have demonstrated significant promise in predicting household-level wealth indices, enabling the creation of high-resolution wealth maps that can be leveraged across multiple causal trials while addressing chronic data scarcity in global development research. However, because standard training objectives prioritize overall predictive accuracy, these predictions often suffer from shrinkage toward the mean, leading to attenuated estimates of causal treatment effects and limiting their utility in policy evaluations. Existing debiasing methods, such as Prediction-Powered Inference (PPI), can handle this attenuation bias but require additional fresh ground-truth data at the downstream stage of causal inference, which restricts their applicability in data-scarce environments. We introduce and evaluate two post-hoc correction methods—Linear Calibration Correction (LCC) and a Tweedie’s correction approach—that substantially reduce shrinkage-induced prediction bias without relying on newly collected labeled data. LCC applies a simple linear transformation estimated on a held-out calibration split; Tweedie’s method locally de-shrink predictions using density score estimates and a noise scale learned upstream. We provide practical diagnostics for when a correction is warranted and discuss practical limitations. Across analytical results, simulations, and experiments with Demographic and Health Surveys (DHS) data, both approaches reduce attenuation; Tweedie’s correction yields nearly unbiased treatment-effect estimates, enabling a “one map, many trials” paradigm. Although we demonstrate on EO-ML wealth mapping, the methods are not geospatial-specific: they apply to any setting where imputed outcomes are reused downstream (e.g., pollution indices, population density, or LLM-derived indicators).
@@ Introduction
Machine learning (ML) models trained on Earth observation (EO) data have recently demonstrated impressive performance in predicting household-level wealth indices, such as the International Wealth Index (IWI), with reported $R^{2}$ values approaching 0.80 (pettersson2023time; jean2016combining; daoud2023using; chi2022microestimates). These EO-ML models offer a much-needed solution to longstanding data scarcity issues in global development, enabling researchers to generate high-resolution, retrospective “wealth maps” in data-poor regions. Indeed, two recurring downstream uses are: (i) evaluating aid or infrastructure programs (e.g., differences in post-intervention wealth across treated vs. control regions), and (ii) tracking regional poverty trends over time.
But while these predictions are promising in the EO-ML upstream task, a key question remains (jerzak2023integrating; jerzak2023image): the value of EO-ML-derived data hinges on whether domain researchers, in the downstream phase, can use these data to estimate causal effects for their specific application areas, from evaluating applied environmental to social science questions (daoudStatisticalModelingThree2023). However, even under ideal conditions, the predictions of EO-ML models tend to have lower variance than the target values, leading to attenuated treatment effects and, thus, potentially misleading conclusions (berglund2012regression). When upstream predictions shrink toward the global mean, estimated effects are biased toward zero, increasing Type-II error: a true 5% effect can appear as only 2–3%, inviting “no impact” conclusions despite meaningful gains. Figure 1 provides a pipeline overview linking the upstream map to multiple downstream trials without requiring new labels.
@@
Recent research has developed methods that can debias upstream machine-learning model predictions, but they require either fresh data in the downstream stage (olofsson2013making; dsl_original; lu2024quantifying) or changing the way the model is trained (ratledge_using_2022). For example, relying on such fresh data, angelopoulos2023prediction propose Prediction-Powered Inference (PPI), a method that corrects any ML model for downstream inference. Similarly, dsl_original achieve the same goal but with better sample efficiency (depieuchon2025benchmarkingdebiasingmethodsllmbased), but it also requires knowing the probability of observing a sample point. While these methods offer a path to integrate upstream and downstream analyses, their success depends on the availability of newly collected treatment-predictor-outcome tuples in the downstream causal step or knowledge of the data labeling process (dsl_original). Often, such data are prohibitively costly or unethical to collect.
@@
ratledge_using_2022 address the challenge of attenuation bias and the unavailability of data in the downstream by adjusting the model’s loss function instead. During training, they penalize quintile-level discrepancies between observed and predicted means, encouraging predictions to match the true outcome distribution more closely across its support. Although their solutions do not need additional downstream data, they require modifying the training pipeline, which is computationally expensive and can degrade predictive performance.
@@
Thus, to enable broader applicability of ML models, the upstream team needs to create a data product that is agnostic to downstream use, requires no downstream-fresh data, produces unbiased causal estimates (due to attenuation), and is computationally efficient (daoudStatisticalModelingThree2023).
@@
We address this gap by proposing a linear calibration method and a family of Tweedie corrections, a novel approach that reduces prediction bias without requiring fresh labeled data. We evaluate these two approaches against three others: Prediction-Powered Inference (PPI), the loss-adjustment approach of ratledge_using_2022, and naive (uncorrected) predictions. The methods introduced here pave the way for the same wealth map to be reused across multiple trials without needing to refit a model for each trial. This also removes the requirement for interaction between the upstream machine-learning team creating the data product and any other domain-knowledge team in the downstream phase. Hence the concept: one map, many trials. We derive analytical proofs and demonstrate the empirical performance of all methods in both simulations and real-world applications, using data from the African Demographic and Health Survey (DHS).
We make these corrections accessible via an open-source software package called unshrink (see Appendix VI); although developed for EO-ML wealth maps, these same attenuation and remedies arise broadly whenever surrogate outcomes replace direct measurements, including in environmental (air pollution, urban heat), demographic (population density), and text/LLM-annotated constructs.
Figure 1: Pipeline for causal inference using ML predictions without fresh ground-truth data. Upstream teams train and calibrate once; downstream teams reuse the same map across multiple trials.
Problem Setup & Related Work
@@
Statistical predictions can systematically shrink estimates toward the central modes of the training distribution (ting2024machine). This phenomenon is rooted in optimal error-minimization estimation (cf. Stein’s paradox), where a model “hedges” predictions in regions of greater uncertainty. When facing noise or limited training data, the optimal strategy, in terms of minimizing residuals, is to pull predictions toward denser regions of the training data. For instance, consider an image that appears with an IWI of 80, but the model has observed this outcome only once, compared to thousands of observations with an IWI of around 70. A perfect prediction would be 80 for this data point, but the model will pull this data point’s prediction towards 70 instead to hedge against the risk of overcommitting to a rare value.
@@
When such biased predictions replace true outcomes in downstream causal analyses, the resulting regression-to-the-mean induces attenuation bias, reducing observed treatment group differences even under randomization (shu2019causal). The difference in mean predicted outcomes between treated and control units shrinks both group means toward the global average, thereby attenuating the estimated effect. In our held-out DHS clusters, the calibration curve exhibits this shrinkage pattern—overpredicting poor locations and underpredicting wealthy ones (Figure 2).
@@
Regression-to-the-mean bias is particularly acute in upstream tasks with scarce ground truth (due to finite samples) or in high-dimensional input data (due to the curse of dimensionality), forcing greater reliance on error-prone predictions—something especially pertinent in development economics, political science, and global health contexts—where household data is often expensive or unavailable—to increasingly substitute upstream predictions (kino2021scoping; kakooei2024mapping; daoudImpactAusterityChildren2024). Moreover, high out-of-sample $R^{2}$ does not eliminate this bias: models can achieve strong accuracy yet yield attenuated causal estimates (ting2024machinelearningmodelssystematically).
@@
Existing remedies are PPI, which subtracts a rectifier estimated from fresh ground-truth pairs at the causal-inference stage (angelopoulos2023prediction), or loss-function adjustments that penalize quintile-level errors (ratledge_using_2022). However, these approaches require that the upstream team adapts the data for downstream analysis, slowing or limiting altogether “one map, many trials” efforts.
@@
To address these limitations, we propose a Tweedie-inspired approach. This Tweedie approach debiases the shrinkage effect in ML-predicted outcomes, enabling downstream researchers to accurately recover causal effects from the upstream data product. Our Tweedie framework treats the ML model as a black box: we assume access only to its out‑of‑sample predictions and do not rely on retraining or modifying the model. Under this setting, we demonstrate that shrinkage can be reversed, either by learning a global calibration correction from held-out labeled data or by estimating a local correction using ideas from Tweedie’s formula, initially described by Robbins and Tweedie in 1956 in the context of empirical Bayes estimation of means under Gaussian observation noise. Both strategies aim to recover unbiased treatment effect estimates from biased predictions, without requiring fresh ground-truth labels at the time of intervention. We benchmark our Tweedie adaptation to existing methods, which are compared in Table A.III.4 based on their associated assumptions and data/training requirements. [^1] As previously mentioned, the only method we are unable to deploy in our setting is the DSL method, as the probability of DHS samples depends on unknown country-level DHS sampling strategies (probabilities not publicly provided at the unit level). In short, Tweedie’s formula is canonical; we adapt it to post-hoc, local de-shrinkage of surrogate outcomes used for causal estimation without new labels, estimating the density score and noise scale from upstream data and applying correction downstream at analysis time.
@@
Setup. Our pipeline assumes access to two datasets. First, from upstream, a labeled dataset with training points (with indices in $\mathcal{I}_{\textrm{Train}}$): $D_{\text{Train}}=\{(\mathbf{X}_{i},Y_{i}):i\in\mathcal{I}_{\textrm{Train}}\},$ where $\mathbf{X}_{i}\in\mathcal{X}$ denotes covariates (e.g., satellite imagery) and $Y_{i}$ is ground-truth outcome (e.g., poverty level from the DHS).
@@
Next, in the downstream phase, we assume that different teams have assembled collections of trial datasets indexed by $s$, corresponding to different randomized controlled trials, quasi-experiments, or geographic partitions, with a specific treatment (or action) of interest indexed by treatment setting: $D_{\text{Trial $s$}}=\{(\mathbf{X}_{i},\hat{Y}_{i},A_{is}):i\in\mathcal{I}_{\textrm{Trial $s$}},s\in\mathcal{S}\}.$ The data product associated with the outcome is population-wide (e.g., all-Africa), but investigators may want to use that data in multiple causal settings, $s$, each with a different treatment, $A_{is}\in\{0,1\}$. The outcome $Y_{i}$ is not directly observed for all ii. Instead, inference will need to take place through the noisy variable, $\hat{Y}_{i}$. We henceforth remove dependence on $s$ for simplicity.
@@
A key idea here is that the upstream team—a machine-learning team—creates an agnostic dataset (grimmer2022text), meaning that this team’s data product should produce unbiased causal estimates for any number of applications (trials) for a variety of (social science) teams downstream. This setup implies that the upstream team versus the downstream teams should be able to execute on their respective research goals without communication—hence the firewall between the two phases.
@@
In this section, we present a suite of methods designed to debias treatment effect estimates that rely on ML-predicted outcomes. Our overarching goal is to mitigate regression-to-the-mean dynamics in the downstream analysis, improving estimates of causal effects or population-level statistics. We first more formally describe two existing solutions, PPI and the method of ratledge_using_2022, before introducing two no-new-data debiasing approaches—(1) Linear Calibration Correction (LCC) and (2) Tweedie’s Corrections—that aim to correct shrinkage bias without requiring fresh labels in the causal step.
Existing debiasing methods (PPI, Ratledge, DSL)
@@
PPI proposes a framework for valid statistical inference when a large set of unlabeled observations is paired with a predictive model’s outputs (angelopoulos2023prediction). Concretely, letting $\theta^{*}$ be the target parameter (e.g., a population mean) and $f$ denote the model that maps covariates $\mathbf{X}_{i}$ to predicted outcomes $f(\mathbf{X}_{i})$, PPI forms an estimator by adding a rectifier term—an adjustment for the average discrepancy between $Y_{i}$ and $f(\mathbf{X}_{i})$ observed in a smaller, labeled dataset. An estimate for $\theta^{*}$ is computed using an “imputed” version based on unlabeled data ($\{\mathbf{X}_{i}\}_{i={n+1}}^{N}$), then corrected by subtracting the bias estimated on the labeled subset. For mean estimates, this adjustment is:
$\hat{\theta}^{\textrm{PPI}}=\frac{1}{N-n}\sum_{i={n+1}}^{N}f(\mathbf{X}_{i})-\frac{1}{n}\sum_{i=1}^{n}\left(f(\mathbf{X}_{i})-Y_{i}\right)$
@@
Because the rectifier is estimated on ground-truth pairs $\{\mathbf{X}_{i},Y_{i}\}_{i=1}^{n}$, the procedure maintains valid coverage for $\theta^{*}$ in large samples, provided that the predictive model $f$ is fixed (i.e., not retrained on the labeled data). Unlike our approach, however, PPI crucially relies on gathering labeled observations in the downstream phase. Such labels come from high-quality surveys coordinated by USAID across the African continent and cost millions of US dollars to execute. Given the 2025 funding cuts to USAID, it is unlikely that new and fresh DHS data points will be collected. Hence, there is an acute need for unbiased data creation, which should be fully handled in the upstream phase. We will nonetheless include PPI correction in our analysis with the goal of showing what would have happened to our estimates had fresh data actually been available downstream.
@@
Another approach to remedy attenuation bias is to adjust the loss function: ratledge_using_2022 propose a specialized upstream training procedure that directly penalizes quintile-level discrepancies between observed and predicted means, thereby encouraging $\hat{Y}_{i}$ to match the true distribution of $Y_{i}$ more closely across the support of the outcome distribution. Specifically, the training objective augments the mean squared error loss with an $\ell_{2}$ penalty and a quintile‑bias term:
$\mathcal{L}(\theta)\;=\;\mathrm{MSE}(\theta)\;+\;\lambda_{r}\,\|\theta\|_{2}^{2}\;+\;\lambda_{b}\,\hat{E}_{b}(\theta),$
where MSE($\theta$) is the mean squared error given parameters, $\theta$, $\lambda_{r}\,\|\theta\|_{2}^{2}$ is a regularization term, and $\lambda_{b}\hat{E}_{b}(\theta)$ is the quintile-specific bias loss, with $\hat{E}_{b}(\theta)=\max_{j}\{\hat{B}_{j}(\hat{f}_{\theta})^{2}\}$ and $\hat{B}_{j}(\hat{f}_{\theta})=\mathbb{E}[\hat{f}_{\theta}(\mathbf{X}_{i})-Y_{i}\mid Y_{i}\in Q_{j}]$ penalizing quintile‑level discrepancies (where quantiles are indexed by $j$). Here, $\lambda_{r}$ and $\lambda_{b}$ denote the weighting terms for the regularization and quantile-specific loss, respectively. [^2] In our experiments, we use AdamW with weight decay for $\ell_{2}$ regularization; to avoid double‑counting, we set $\lambda_{r}=0$ whenever weight decay is non‑zero.
@@
Through the quintile-specific bias loss, this approach reduces the likelihood that group-level estimates of causal effects are attenuated when $\hat{Y}_{i}$ is used in downstream analyses by directly operating on model parameters, $\theta$. In contrast, the two new methods introduced next operate as post-hoc corrections on out-of-sample predictions from a black-box model, requiring no specialized retraining or customized loss function that might reduce predictive performance.
1. Linear Calibration Correction (LCC)